OpenStudy (anonymous):

please medal me i give you 100 dollar bill

OpenStudy (loser66):

hehehe....

OpenStudy (anonymous):

100 dollar bill that you have to p;ay

OpenStudy (anonymous):

its a joke

OpenStudy (loser66):

The night is upcoming. Post the question; I will help if I can, then I sleep and dream I get 100 dollars from you. hehehe

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

(1 pt) Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85

OpenStudy (anonymous):

i think d out of these questions A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Yes, No need to wait to night; I am day dreaming now. I see someone tells me that the last one is the answer. hehehe

OpenStudy (anonymous):

ok i medal you you medal me

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Let me wait until night, then I will give you medal in dream. In day, you do nothing to get medal.

OpenStudy (anonymous):

yes i did what ever